OpenStudy (anonymous):

Jacob spends 60 minutes in the gym every day doing freehand exercises and running on the treadmill. He spends 30 minutes more running on the treadmill than doing freehand exercises. Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does freehand exercises (y) and the number of minutes he runs on the treadmill(x). (5 points) Part B: How much time does Jacob spend on doing freehand exercises? (3 points) Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points)

OpenStudy (anonymous):

I need help w/ Part C?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Here are my answers: Part A.) A pair of linear equations would be: x+ y = 60, this would be because he spent 60 minutes doing both of his exercises. x = y + 30 Part B.) He spent 15 minutes doing freehand exercises.

OpenStudy (phi):

Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points) can x= 40 ? we have x=y+30 or 40= y+30 --> y= 10 minutes doing freehand x+y= 40+10 = 50 minutes He stays at the gym for 60 minutes. It is possible he goofs off for 10 minutes, and does exercises for 50 minutes. So I would say yes, he could have done 40 minutes of running.

OpenStudy (anonymous):

awesome, thank you phi, you're my hero

OpenStudy (texaschic101):

y = freehand and x = treadmill y + x = 60 x = y + 30 y + y + 30 = 60 2y = 60 - 30 2y = 30 y = 15 freehand is 15 minutes x + y = 60 x + 15 = 60 x = 60 - 15 x = 45 yes....he can spend up to 45 minutes doing treadmill

OpenStudy (anonymous):

thank you both

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

If I need more help, may I ask you more? @texaschic101

OpenStudy (texaschic101):

you can ask....I can't make no promises that I can help, but I will try

OpenStudy (anonymous):

Ok, here's another question: A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@texaschic101 did I anger you? :(

OpenStudy (texaschic101):

no you did not anger me...why would you think that ? I am sorry but I am not sure how to do these types of problems.

OpenStudy (anonymous):

ok